OpenStudy (anonymous):

Who knows how to do the distance formula? I need help. It always confuses me!! Find the length of the line segment whose endpoints are (- 5, 6) and (6, 6).

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 6}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ 6}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (mathstudent55):

The distance between points \((x_1, y_1) \) and \((x_2, y_2) \) is \(d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \) So far so good?

OpenStudy (mathstudent55):

Where do you get confused?